# Training Small Multimodal Models to Bridge Biomedical Competency Gap: A Case Study in Radiology Imaging

### 12 Mar 2024 v1

@@ Methods
F1-14 F1-5 F1-14 F1-5 F1-14 F1-5 F1-14 F1-5 F1 (1) (4) -L
LLaVA-Rad 7B 57.13 57.32 41.37 48.51 56.73 59.69 44.44 52.29 29.59 38.61 15.48 30.80
Med-PaLM M 84B 53.56 57.88 39.83 51.60 - - - - 26.71 32.31 11.31 27.29
GPT-4V - 35.50 25.83 20.39 19.64 35.56 33.31 25.27 29.63 13.21 16.38 1.91 13.22
MAIRA-1 7B 55.70 56.00 38.60 47.70 55.30 58.80 42.30 51.70 24.30 39.20 14.20 28.90
LLaVA-Med 7B 27.22 22.01 15.48 16.55 27.26 24.40 18.72 20.53 6.49 22.23 1.01 13.29
LLaVA 7B 22.92 23.39 15.38 17.48 23.74 26.90 17.02 20.31 4.47 21.02 1.29 13.83
CheXagent 7B 25.61 18.70 15.37 15.63 25.90 19.77 17.64 16.94 17.97 14.34 3.88 20.86
Flamingo-CXR <1B - - - - 51.90 56.50 - - 20.50 - 10.10 29.70
CvT2Dist. <1B 44.20 - 30.70 - - - - - - 39.28 12.67 28.63
$\mathcal{M}^{2}$ trans <1B - - - - - 56.70 - - - - 11.40 -
RGRG <1B - - - - - 54.70 - - - 37.30 12.60 26.40
R2Gen <1B - - - - 22.80 34.60 - - - 35.30 8.60 27.70
TieNet <1B - - - - - 27.10 - - - - 8.10 -
@@
LLaVA-Rad 57.13 57.32 41.37 48.51 56.73 59.69 44.44 52.29 29.59 38.61 15.48 30.80
Variant #1 55.47 55.21 38.36 46.92 55.87 58.35 42.50 51.14 29.60 38.46 15.38 30.56
Variant #2 55.57 56.42 38.50 48.55 55.65 59.45 41.62 52.94 28.94 37.47 15.16 30.32
Variant #3 55.68 56.20 39.14 47.39 55.80 59.04 43.08 51.43 29.70 37.86 15.63 30.98
Variant #4 56.15 56.89 38.55 48.26 56.09 60.19 43.03 54.04 27.62 29.85 10.02 26.30
Variant #5 50.62 53.27 30.70 42.74 49.19 54.09 33.59 45.00 25.02 16.89 4.97 16.36
Variant #6 51.11 52.87 31.00 43.59 51.41 56.26 35.10 48.58 26.88 34.95 13.93 29.44
Variant #7 49.75 52.05 30.48 42.44 50.27 55.41 34.47 47.97 26.53 34.44 13.73 29.12
Variant #8 49.62 52.18 31.34 42.83 50.07 55.56 35.02 47.79 26.90 34.27 13.69 29.26
Analysis #1 96.71 98.37 95.06 97.82 96.88 98.39 95.50 97.78 93.19 82.19 78.82 90.88

### 20 Mar 2024 v2

@@ Methods
predictions, wts = weights F1-14 F1-5 F1-14 F1-5 F1-14 F1-5 F1-14 F1-5 F1 (1) (4) -L
LLaVA-Rad 7B 57.1 57.3 41.4 48.5 56.7 59.7 44.4 52.3 29.6 38.6 15.5 30.8
Med-PaLM M 84B 53.56 57.9 39.8 51.6 - - - - 26.7 32.3 11.3 27.3
GPT-4V - 35.5 25.8 20.4 19.6 35.6 33.3 25.3 29.6 13.2 16.4 1.9 13.2
MAIRA-1 7B 55.7 56.0 38.6 47.7 55.3 58.8 42.3 51.7 29.6 39.2 14.2 28.9
LLaVA-Med 7B 27.2 22.0 15.5 16.6 27.3 24.4 18.7 20.5 6.5 22.2 1.0 13.3
LLaVA 7B 22.9 23.4 15.4 17.5 23.7 26.9 17.0 20.3 4.5 21.0 1.3 13.8
CheXagent 7B 25.6 18.7 15.4 15.6 25.9 19.8 17.6 16.9 18.0 14.4 3.9 20.9
Flamingo-CXR <1B - - - - 51.9 56.5 - - 20.5 - 10.1 29.7
CvT2Dist. <1B 44.20 - 30.70 - - - - - - 39.3 12.7 28.6
$\mathcal{M}^{2}$ trans <1B - - - - - 56.7 - - - - 11.4 -
RGRG <1B - - - - - 54.7 - - - 37.3 12.6 26.4
R2Gen <1B - - - - 22.8 34.6 - - - 35.3 8.6 27.7
TieNet <1B - - - - - 27.1 - - - - 8.1 -
@@
LLaVA-Rad 57.1 57.3 41.4 48.5 56.7 59.7 44.4 52.3 29.6 38.6 15.5 30.8
Variant #1 55.5 55.2 38.4 46.9 55.9 58.4 42.5 51.1 29.6 38.5 15.4 30.6
Variant #2 55.6 56.4 38.5 48.6 55.7 59.5 41.6 52.9 28.9 37.5 15.2 30.3
Variant #3 55.7 56.2 39.1 47.4 55.8 59.0 43.1 51.4 29.7 37.9 15.6 31.0
Variant #4 56.2 56.9 38.6 48.3 56.1 60.2 43.0 54.0 27.6 29.9 10.0 26.3
Variant #5 50.6 53.3 30.7 42.7 49.2 54.1 33.6 45.0 25.0 16.9 5.0 16.4
Variant #6 51.1 52.9 31.0 43.6 51.4 56.3 35.1 48.6 26.9 35.0 13.9 29.4
Variant #7 49.8 52.1 30.5 42.4 50.3 55.4 34.5 48.0 26.5 34.4 13.7 29.1
Variant #8 49.6 52.2 31.3 42.8 50.1 55.6 35.0 47.8 26.9 34.3 13.7 29.3
Analysis #1 96.7 98.3 95.1 97.8 96.9 98.4 95.5 97.8 93.2 82.2 78.8 90.9